5x^-8/4^-3z^2


5x^-8/z^2/4^3
5x^-8/z^2/64
320/x^8z^2
is there a break down you could show me?

Answer by richwmiller(17219)

(5x^(-8))/(4^(-3)*z^2) does equal 320/(x^8 z^2)
Remember
5x^(-8)=1/5x^8
and
4^(-3)=1/4^3
so you flip
5*4^3/x^8*z^2
5*64/x^8*z^2
320/(x^8 z^2)